Fermi National Laboratory
 
Operation Summary for February 8, 2012 - February 10, 2012
 

Wednesday February 8

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam, but with reduced intensity to NuMI due to vacuum quality after the MI magnet replacement.

During the day shift, at 8:19 AM, the Linac beam permit tripped off due to bad beam positions in the 400 MeV area; it reset. Interlock personnel began conducting an hour-long safety system test of the MTest enclosures at 8:36 AM. A Linac Klystron RF station (KRF3) tripped off at 8:36 AM; it reset. At 10:04 AM, Operations attempted to send beam to the next FTBF Meson experiment T-1017, but problems with two power supplies (LAM52 and V701) tripping off stopped them; Operators investigated and found that the supplies were stuck in 8 GeV mode. Operators fixed the problem and Operations began sending beam to T-1017 at 10:26 AM. At 11 AM, a Lambertson magnet power supply (LAM10) started tripping off; Power Supply personnel investigated and found a bad SCR firing card and a bad soft start resistor, which held off beam for about an hour and twenty minutes during repairs. Operations turned on the Muon Rings power supplies at 11:12 AM. Water Group personnel reported at 1:48 PM that they were valving in the MI-8 LCW system. A Muon Rings person began a two and half hour study at 2:47 PM. At 3:25 PM, KRF1 tripped off; it reset.

During the evening shift, Operators reported at 6:23 PM that they added water to the Meson MS2 LCW system. At 6:50 PM, Operators reported having trouble keeping stable beam intensity to T-1017; an Operator made a small change to a power supply that increased efficiency. After fixing some small problems, Operators injected protons into the Recycler at 9:24 PM. Operations disabled beam to T-1017 at 10 PM. Operators reported at 11:27 PM that there had been several NuMI beam permit trips this shift due to incomplete RF gaps, but no RF stations had tripped off; everything reset. At 11:38 PM, MIRF8 tripped off; it reset.

Thursday February 9

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 1:06 AM, Operators reported problems with the 750 KeV area Buncher causing beam to jump up and down; Linac personnel came in and worked on the problem with the Linac low energy Buncher. Linac resumed sending beam at 3:16 AM. Operators began making changes to the MI transition time at 4:51 AM. Operators reported at 6:12 AM that they were still having problems with the Linac Buncher gradient; Operators swapped the gradient regulator waveform generator. At 7:44 AM, the Neutron Therapy Facility (NTF) personnel prepared to treat two patients later in the morning.

During the day shift, at 8:28 AM, a Water Group person reported that CUB Column #1, which served MI, was clogged and needed cleaning. At 9:16 AM, LRF2 tripped off; it reset. Operations established beam to T-1017 around 10:20 AM. At 12:03 PM, LRF3 tripped off for the first of three times in 15 minutes; it reset. At 3:26 PM, KRF3 tripped off; it reset. At 3:35 PM, BRF2 tripped off; it reset. Water Group personnel reported that Column #1 was back online.

During the evening shift, at 4:50 PM, LRF4 tripped off; it reset. At 4:58 PM, BRF11 tripped off; it reset. At 5:14 PM, KRF1 tripped off for the first of two times in 40 minutes; it didn’t reset easily after the first trip and a Linac technician had to bring it back after the second trip. At 5:24 PM, a Linac quadrupole magnet power supply tripped off; it reset. Operations disabled the T-1017 beam a little after 6 PM; later, the experiment reported that a key piece of equipment had failed and they would have to end their current run. At 9:05 PM, LRF3 tripped off; it reset. Operators reported problems with Booster beam positions they couldn’t resolve; a Booster person came in and found that a new phase detector they had recently installed was the cause. He made a slight adjustment that resolved the problem. At 11:19 PM, KRF5 tripped off; it reset. At 11:49 PM, a Booster vacuum pump (IP22) tripped off and took down BRF11 and 12; everything recovered.

Friday February 10

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 1:01 AM, LRF4 tripped off; it reset. Operators reported more problems with beam positions in Booster; they adjusted the Phase Detector as previously instructed. Operators reported that they had tuned the MI dampers at 3:59 AM to improve efficiencies. At 4:13 AM, KRF3 tripped off; it took multiple resets to get it back on. At 5:09 AM, KRF3 tripped off again; Operators found the +15-volt power supply light indicator out and a blown fuse for the PFN control module. Operators replaced the fuse, saw the +15-volt light appear for a moment and then go out again, but the RF station stayed on. At 5:25 AM, KRF3 tripped off again; operators replaced the fuse, but the fuse blew out immediately. An Operator called in a Linac tech. At 5:35 AM, BRF11 tripped off; it reset. Operations switched to an energy conservation time line at 5:54 AM. At the same time, Operations received a LCW leak alarm. Operators prepared enclosures C, D, E, and F for access. The H- Source high voltage tripped off at 7:19 AM; it reset. At 7:40 AM, NTF personnel prepared to treat two patients later this afternoon.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with plans by Booster, MI, Muon Rings, and SY personnel to conduct studies.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/10/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory